# Automatic Model Extraction of the Match Standard in Symmetric–Reciprocal–Match Calibration

### v1

@@ Abstract
This paper addresses the modeling of parasitics of the match standard in the symmetric-reciprocal-match (SRM) calibration method of vector network analyzers (VNAs). In the general SRM procedure, the match standard is assumed to be fully known. Here, we demonstrate that the match can be modeled with an arbitrary frequency-dependent model using a non-linear global optimization procedure. To highlight the validity of the suggested approach, numerical tests were conducted, demonstrating the ability to recover the match standard parasitic model down to software numerical precision. Additionally, we performed microstrip line measurements to compare the SRM calibration with match modeling to the multiline thru-reflect-line (TRL) calibration one, showing that automatic model extraction can achieve accuracy similar to using a match standard defined through multiline TRL calibration.
@@ I Introduction
The common calibration procedures for vector network analyzers (VNAs) rely on known standards. Most common procedure is the short-open-load-thru (SOLT) method [1], or with an unknown reciprocal device in the short-open-load-reciprocal (SOLR) method [2]. VNA calibration has evolved and improved in accuracy by introducing new methods that rely on partially defined standards. Such calibration techniques are summarized as self-calibration procedures [3]. Well-known and often applied representatives of self-calibration procedures are, among others, the aforementioned SOLR method, the thru-reflect-line (TRL) and the multiline TRL [4, 5, 6, 7], the line-reflect-match (LRM), the thru-match-reflect-reflect (TMRR), and the line-reflect-reflect-match (LRRM) methods [8, 9, 10, 11].
The symmetric-reciprocal-match (SRM) method was recently introduced in [12], which relies solely on the symmetry and reciprocity properties of the standards while requiring a fully defined match standard. This method offers advantages over conventional self-calibration procedures, such as LRM/TMRR/LRRM, as it eliminates the need for a defined thru/line during calibration.
This paper extends the SRM method by incorporating automatic parasitic extraction modeling of the match standard. The enhancement requires only knowledge of the match standard’s DC resistance, while its parasitic response is modeled as a frequency-dependent equivalent circuit. Unlike the LRRM method [11], which uses a simple inductance model, the SRM approach allows for more complex match standard structures through global optimization.
A significant advantage of the proposed method is that all standards are partially defined, except for the match standard, whose DC value is needed to establish the reference impedance definition of the calibrated S-parameters.
@@
The article is organized as follows. Section II reviews the general procedure for the SRM method. Section III discusses the parasitic modeling of the match and the optimization procedure. Section IV presents numerical analysis using synthetic data and experimental measurements on printed circuit boards (PCB), followed by conclusions in Section V.
@@ II SRM Calibration Review
The reason for expressing the one-port measurements as Möbius transformations is that the operations can be composed, which translates to a simple matrix product for the coefficients [14]. This allows us to easily express the inverse of the operation by taking the inverse of the matrix of the coefficients. For example, for the port-B measurement from (2b), the input coefficient can be reverse calculated as follows:
@@
Therefore, with $\boldsymbol{H}$ and $\boldsymbol{F}$ and the measurement of the unknown transmissive device $\boldsymbol{M}_{\mathrm{net}}$ from (1), it can be shown that a “virtual” thru standard can be created that can be used to formulate an eigenvalue problem for both port-A and port-B coefficients. Reference [12] discusses this in greater detail, and provides the equations for the case of measuring the cascaded one-port standards at port-B.
@@
The system of equations for port-B can be obtained in a similar way, resulting in the following system of equations:
@@ III Automatic Parasitic Model Extraction
In the derivation outlined in Section II, (9) and (10) were obtained, which can only be solved if a fully defined match standard is used. In many cases, the match standard can be challenging to accurately define at higher frequencies, as parasitic elements become more impactful as the wavelength approaches the device’s physical size. This is further demonstrated with measurements in Section IV.
@@
While the equivalent circuit model presented in Fig. 2 is an example of a match standard, it could also be used to describe non-ideal open or short standards, where $\rho_{\mathrm{dc}}=1$ is for an open standard and $\rho_{\mathrm{dc}}=-1$ is for a short standard. However, in general, the complexity of the model depends on how the match standard is implemented in reality. For example, in the measurement example discussed in Section IV-B, a flip-chip resistor is used, which needed to be described with a more complex lumped model to account for the mounting pads and the resistor’s internal parasitics.
The question is: if a one-port device can be described using an equivalent circuit model, how can the unknown parameters be solved without the match standard’s measurement? This can be addressed through non-linear optimization of the nullspace problem through singular value decomposition (SVD). Recalling the system of equations earlier in either (9) or (10), this system can be expanded by including other one-port standards besides the match standard, as shown in (11) below for port-A:
@@
where $f_{i}$ refers to the $i$th frequency point, and $N$ is the total number of frequencies. Once $\boldsymbol{\theta}_{\mathrm{opt}}$ have been found, the parameters associated with the match standard are used to define its reflection coefficient accurately.
@@
After $\boldsymbol{\theta}_{\mathrm{opt}}$ is found, the coefficients associated with the match standard are substituted back into (9) and (10). If using symmetric match pairs, optimization can be performed on only one side. However, for asymmetric match standards, each port must be solved independently. Furthermore, while additional unknown standards can be included in the SVD procedure (e.g., both short and open standards), this significantly increases the number of parameters to be solved, as the unknown parameters of each standard are not coupled. This makes it challenging even for global optimization methods to find all parameters simultaneously. Therefore, it is best practice to minimize the complexity of the optimization whenever possible, either through number of solved parameters or coupling the parameters in both ports. Additionally, the speed of the optimization is also a function of the evaluation of the models. For example, incorporating a transmission line segment, as in the numerical example in Section IV-A, can be computationally expensive when using special integral functions as the Bessel functions [17].
@@ IV-A Numerical Analysis
The numerical analysis involves creating synthetic data of co-planar waveguide (CPW) standards using the model developed in [17, 18, 19]. To accurately simulate an on-wafer setup, error boxes extracted from actual on-wafer measurements are utilized using multiline TRL calibration on a CPW impedance substrate standard (ISS). For detailed information about the measurement setup, calibration substrate, and measurements, readers should refer to [7]. The goal is to generate SRM standards based on the CPW model and embed them within the error boxes of an actual VNA setup. Fig. 3 illustrates the experimental setup.
@@
The CPW cross-section structure has the following dimensions: signal width of $49.1\,\mu\mathrm{m}$, ground width of $273.3\,\mu\mathrm{m}$, conductor spacing of $25.5\,\mu\mathrm{m}$, and conductor thickness of $4.9\,\mu\mathrm{m}$. The substrate is Alumina with a real-part relative permittivity of $\epsilon_{r}^{\prime}=9.9$ and loss tangent of $\tan\delta=0.0002$. The conductor is gold with relative conductivity to copper of $\sigma_{r}=0.7$, where copper conductivity is $\sigma=58\,\mathrm{MS/m}$.
The one-port SRM standards were implemented as non-ideal match, short, and open standards with a $200\,\mu\mathrm{m}$ offset, as shown in Fig. 4. For the network-load standards, a $4\,\mathrm{mm}$ line is used as the reciprocal standard, combined with the non-ideal match, short, and open standards. Note that the CPW has a complex-valued, frequency-dependent characteristic impedance that deviates slightly from $50\,\Omega$. Thus, the CPW line segments introduce both an offset and impedance mismatch, as $Z_{\mathrm{ref}}=50\,\Omega$ is used as the reference impedance for the lumped elements. The device under test (DUT) used for verification is a stepped impedance using the same CPW structure but with a $15\,\mu\mathrm{m}$ signal width.
Figure 4: Models used to simulate non-ideal load standards (a) $50\,\Omega$ match standard with $L_{0}=25\,\mathrm{pH},C_{0}=1\,\mathrm{fF}$, (b) short standard with $L_{0}=30\,\mathrm{pH},L_{1}=10\times 10^{-12}\,\mathrm{pH/Hz},C_{0}=0.5\,\mathrm{fF}$, and open standard with $C_{0}=15\,\mathrm{fF},C_{1}=5\times 10^{-15}\,\mathrm{fF/Hz},L_{0}=1\,\mathrm{pH}$. All standards include a $200\,\mu\mathrm{m}$ CPW line offset. These models are used for automatic parasitic extraction.
Only the match and short standards have been used for automatic model parameters estimation. All lumped elements and material properties ($\epsilon_{r}^{\prime}$, $\tan\delta$, and $\sigma_{r}$) were assumed unknown. While including the open standard would have been possible, it would have increased the parameter space from 11 parameters (five from the match standard and six from the short standard) to 17 parameters, significantly increasing optimization time. Symmetric models at both ports were assumed to improve optimization efficiency, as asymmetric models would double the parameter count. This consideration is particularly important since the CPW model, though analytical, uses computationally intensive Bessel functions. In the measurement example in Section IV-B, devices have been modeled using only lumped elements without offset, allowing for more parameters and independent port solutions due to faster evaluation.
@@
To validate calibration accuracy, a relative error metric has been defined:
@@
where $x_{\mathrm{est}}$ represents the estimated quantity and $x_{\mathrm{true}}$ is the true value.
Data processing was conducted using Python with the scikit-rf [20] and scipy [21] packages, employing the DE optimization procedure [21]. Fig. 5 shows the calibrated DUT results comparing two cases: using an ideal match standard versus using nonlinear optimization to fit unknown parameters. This ideal match models the match response by a frequency-independent $50\,\Omega$ resistor, even if the match for the simulation is calculated based on the model introduced in Fig. 4, a). The assumption of frequency independence of the match is well applicable for frequencies below $30\,\mathrm{GHz}$. Above this frequency, the ideal match standard definition produces significant errors in both $S_{11}$ and $S_{21}$. The optimization approach achieves much better accuracy, limited only by numerical precision. Fig. 6 demonstrates this by showing the relative error of estimated parameters converging to floating-point precision. It should be noted that due to the computationally expensive CPW model, the 1000 iterations shown in Fig. 6 took approximately 25 minutes to complete (using 150 frequency points).
Figure 5: True S-parameters of the stepped impedance line (DUT) compared with SRM calibration results using ideal match definition versus model fitting procedure.
@@ IV-B PCB Measurement
For multiline TRL standards, line lengths of $\{0,0.5,4,5.5,6.5,8.5\}\,\mathrm{mm}$ have been used, where the first line is the thru line. The reflect standard was implemented as an open standard. For SRM standards, symmetric open, short, and match standards and network-load with an offset line segment were implemented. The match standard uses a $50\,\Omega$ flipchip Vishay CH0402-50RGFPA [24], specified up to $50\,\mathrm{GHz}$ with $\pm 2\%$ DC resistance uncertainty. Our DC measurement showed a resistance value of $49\,\Omega$. The small size of the resistor (footprint size 0402) made manual soldering challenging, introducing asymmetry from solder bumps. Fig. 7 shows the measurement system and soldered resistors.
@@
For comparison between SRM and multiline TRL, a stepped impedance microstrip line was used as DUT with $100\,\mu\mathrm{m}$ strip width. For proper comparison, the multiline TRL calibration was renormalized to a constant reference impedance of $50\,\Omega$. Fig. 8 shows the fitting models used for short, open, and match standards in SRM calibration. Both open and short standards were included in the optimization since they only comprise lumped elements and their evaluation is computationally efficient. The fitting was performed independently at each port. The total optimization time for both ports was approximately four minutes, fitting 22 parameters (using 197 frequency points). Only the parameters associated with the match results are used in the final SRM calibration. The resistor model follows the datasheet specifications [24], with the VIA termination modeled as an inductor.
@@
Fig. 9 compares the S-parameters of the calibrated DUT between multiline TRL and SRM. The DUT response from multiline TRL calibration was used as the reference for calculating relative error based on (13). For SRM calibration, several scenarios were examined for defining the match standard: First, assuming ideal zero reflection (typical choice when frequency-dependent response is unknown); second, using our optimization procedure from Section III to fit the models in Fig. 8; and finally, using the measurement of the match standard based on the multiline TRL calibration.
Results in Fig. 9 demonstrate that assuming an ideal match response leads to substantial errors, particularly above $10\,\mathrm{GHz}$ in $S_{21}$ magnitude. In contrast, the optimization approach yields errors comparable to defining the match standard based on multiline TRL. Some discrepancies exist between the DUT response based on SRM and multiline TRL calibrations, especially near $50\,\mathrm{GHz}$. This deviation may arise because the match standard exhibits behavior similar to a short standard near $50\,\mathrm{GHz}$, as evident in Fig. 10, thereby increasing the sensitivity of SRM calibration due to insufficient unique standards. Additionally, the discrepancy could stem from inconsistent resistor soldering, as asymmetry in the standards affects the calibration as investigated in [12]—an error that becomes amplified when the match standard behaves like a short standard near $50\,\mathrm{GHz}$. Nevertheless, achieving results comparable to multiline TRL using manually soldered flip-chip resistors is significant, considering only the DC resistance was specified during the SRM calibration. A better agreement with multiline TRL could be achieved using machine-placed or ultrasonic ball-bonded resistors, where pad parasitics are less influential.
@@
The reflection coefficient of the match standard is presented in Fig. 10, demonstrating strong agreement between the SRM-optimized model and the multiline TRL measurement. A significant advantage of the SRM method is that it requires only the DC resistance value as input, with all other parameters determined through optimization. This approach provides greater flexibility compared to LRRM calibration, as it allows arbitrary complexity in the models for match, short, and open standards, whereas LRRM imposes restrictions on the open and match standard models [11, 25].

### v2

@@ Abstract
This paper addresses the modeling of parasitics of the match standard in the symmetric-reciprocal-match (SRM) calibration method of vector network analyzers (VNAs). In the original SRM procedure, the match standard is assumed to be defined. Here, we demonstrate that the match can be modeled with an arbitrary frequency-dependent model using a non-linear global optimization procedure. To highlight the validity of the suggested approach, numerical tests were conducted, demonstrating the ability to recover the match standard parasitic model down to software numerical precision. Additionally, we performed microstrip line measurements on PCB to compare the SRM calibration using the proposed automatic model extraction against an SRM calibration where the match standard is explicitly defined via multiline thru-reflect-line (TRL) data. The results show that the automatic model extraction can achieve accuracy comparable to that of a data-based match definition, validating the approach as an effective substitute when explicit characterization of the match standard is unavailable.
@@ I Introduction
The calibration of vector network analyzers (VNAs) is an essential step to compensate for the systematic imperfections of the measurement equipment and the measurement setup. The most widely known calibration procedure for VNAs is the short-open-load-thru (SOLT) method [1], which requires all four standards to be fully characterized. The short-open-load-reciprocal (SOLR) method [2] extends SOLT by replacing the thru with any transmissive reciprocal device. However, both methods are fundamentally limited by the accuracy of the standard definitions, which can be challenging to achieve across broad frequency ranges and different measurement environments. To address this, self-calibration procedures emerged that rely on partially defined standards [3].
Among the most accurate self-calibration methods is multiline TRL [4, 5, 6, 7], which depends on distributed circuit models rather than lumped-element approximations, making it traceable and suitable as a primary calibration [8]. However, multiline TRL becomes impractical at lower frequencies due to extremely long line standards and requires variable distances between measurement planes.
For on-wafer applications, the line-reflect-match (LRM), thru-match-reflect-reflect (TMRR), and line-reflect-reflect-match (LRRM) methods [9, 10, 11, 12] are commonly used. The LRRM method [12] introduced series inductance modeling for the match standard parasitics, later extended to include a transmission line section and a shunt capacitance [13]. However, these simplified models cannot accurately represent various resistor implementations, such as via-based shunt resistors or surface-mounted resistors with parallel capacitance, which are scenarios common in device characterization at the printed circuit board (PCB) level. Furthermore, LRM/TMRR/LRRM require defined line standards for reference-plane positioning, making them challenging when measurement contacts, whether probes or connector launchers, are positioned orthogonally or at angles, as bent lines are more challenging to define [14].
Alternative approaches to calibration include the series-resistor method [15, 16, 17], demanding minimal parasitics and high manufacturing precision for on-wafer applications, and Bayesian optimization methods [18, 19] that physically model all standards. The latter approach is computationally intensive and requires modeling every standard, whereas our method limits modeling to the match standard while exploiting symmetry and reciprocity for the remaining standards.
The symmetric-reciprocal-match (SRM) method [20] combines the advantages of LRRM/LRM/TMRR (undefined symmetric standards) with SOLR (unknown transmissive reciprocal device), requiring only a defined match standard. This is particularly advantageous for PCB environments where standard definition inaccuracies are common [21, 22]. In the original publication on SRM calibration [20], the procedure was demonstrated using coaxial measurements. However, it can also be applied to waveguide scenarios, as defining the measurement standards through symmetry and reciprocity can be easily accomplished with waveguide shims and shorts.
This paper extends SRM by incorporating automatic parasitic extraction of the match standard, requiring only its DC resistance while modeling the frequency-dependent parasitic response. Unlike LRRM’s simplified models [12, 13], our approach employs global optimization to accommodate complex structures, which can include effects like via hole parasitics in shunt resistor implementations. While this work focuses primarily on quasi-TEM microstrip structures on PCBs, where diverse parasitics commonly arise from circuit topology, the method can be extended to non-TEM scenarios, such as waveguides. In such a non-TEM case, instead of relying on DC measurement to anchor the definition of the match, waveguide parameter models are optimized [23].
To validate the proposed approach, we conducted numerical tests to verify the optimization procedure’s ability to recover the underlying model parameters, demonstrating that the match standard parasitic model can be extracted down to software numerical precision. Additionally, we performed PCB microstrip line measurements to compare the SRM calibration obtained using the proposed automatic model extraction with an SRM calibration in which the match standard is explicitly defined via multiline TRL characterization. The results demonstrate that automatic model extraction achieves accuracy comparable to using a data-based match definition, validating the approach as an effective alternative when explicit characterization of the match standard is unavailable.
@@
The article is organized as follows. Section II reviews the general procedure for the SRM method. Section III discusses the parasitic modeling of the match and the optimization procedure. Section IV presents numerical analysis using synthetic data and experimental measurements on PCBs, followed by conclusions in Section V.
@@ II SRM Calibration Review
The reason for expressing the one-port measurements as Möbius transformations is that the operations can be composed, which translates to a simple matrix product for the coefficients [25]. This allows us to easily express the inverse of the operation by taking the inverse of the matrix. For example, for the port-B measurement from (2b), the input coefficient can be reverse calculated as follows:
@@
Therefore, with $\boldsymbol{H}$ and $\boldsymbol{F}$ and the measurement of the unknown transmissive device $\boldsymbol{M}_{\mathrm{net}}$ from (1), it can be shown that a “virtual” thru standard can be created that can be used to formulate an eigenvalue problem for both port-A and port-B coefficients. Reference [20] discusses this in greater detail and provides the equations for the case of measuring the cascaded one-port standards at port-B.
@@
The system of equations for port-B can be obtained similarly, resulting in the following system of equations:
@@ III Automatic Parasitic Model Extraction
In the derivation outlined in Section II, (9) and (10) were obtained. These equations can only be solved if a fully defined match standard is used. In many cases, defining the match standard accurately at higher frequencies can be challenging, as parasitic elements become more significant as the wavelength approaches the device’s physical dimensions. This is further demonstrated with measurements in Section IV.
@@
While the equivalent circuit model presented in Fig. 2 is an example of a match standard, it could also be used to describe non-ideal open or short standards, where $\rho_{\mathrm{dc}}=1$ is for an open standard and $\rho_{\mathrm{dc}}=-1$ is for a short standard. However, in general, the model’s complexity depends on the physical design and manufacturing technology of the match standard. For example, in the measurement example discussed in Section IV-B, a flip-chip resistor is used, which needed to be described with a more complex lumped model to account for the mounting pads and the resistor’s internal parasitics.
The question is: if a one-port device can be described by an equivalent circuit model, how can the unknown parameters be solved without the match standard’s measurement? This can be addressed through non-linear optimization of the nullspace problem through singular value decomposition (SVD). Recalling the system of equations earlier in either (9) or (10), this system can be expanded by including other one-port standards besides the match standard, as shown in (11) below for port-A:
@@
where $f_{i}$ refers to the $i$th frequency point, and $N$ is the total number of frequencies. Once $\boldsymbol{\theta}_{\mathrm{opt}}$ has been found, the parameters associated with the match standard are used to define its reflection coefficient accurately.
@@
After $\boldsymbol{\theta}_{\mathrm{opt}}$ is found, the coefficients associated with the match standard are substituted back into (9) and (10). If using symmetric match pairs, optimization can be performed on only one side. However, for asymmetric match standards, each port must be solved independently. Furthermore, while additional unknown standards can be included in the SVD procedure (e.g., both short and open standards), this significantly increases the number of parameters to be solved, as the unknown parameters of each standard are not coupled. This makes it challenging even for global optimization methods to find all parameters simultaneously. Therefore, it is best practice to minimize the complexity of the optimization whenever possible, either through the number of solved parameters or by coupling the parameters across both ports. Additionally, the speed of the optimization is also a function of the evaluation of the models. For example, incorporating a transmission line segment, as in the numerical example in Section IV-A, can be computationally expensive when using special integral functions as the Bessel functions [28].
@@ IV-A Numerical Analysis
The numerical analysis involves creating synthetic data of co-planar waveguide (CPW) standards using the model developed in [28, 29, 30]. To accurately simulate an on-wafer setup, error boxes extracted from actual on-wafer measurements are utilized using multiline TRL calibration on a CPW impedance substrate standard (ISS). For detailed information about the measurement setup, calibration substrate, and measurements, readers are referred to [7]. However, relevant information about the CPW parameters is presented below. The goal is to generate SRM standards based on the CPW model and embed them within the error boxes of an actual VNA setup. Fig. 3 illustrates the experimental setup.
@@
The simulated CPW cross-section structure has the following dimensions, which were taken from [7]: signal width of $49.1\,\mu\mathrm{m}$, ground width of $273.3\,\mu\mathrm{m}$, conductor spacing of $25.5\,\mu\mathrm{m}$, and conductor thickness of $4.9\,\mu\mathrm{m}$. The substrate is Alumina with a real-part relative permittivity of $\epsilon_{r}^{\prime}=9.9$ and loss tangent of $\tan\delta=0.0002$. The conductor is gold with relative conductivity to copper of $\sigma_{r}=0.7$, where copper conductivity is $\sigma=58\,\mathrm{MS/m}$. It should be noted that the Alumina substrate considered in [7] was assumed to have a loss tangent of zero, as the manufacturer does not provide a value and states it is lossless. We assumed a small $\tan\delta$ of 0.0002 for the simulation, as almost all materials exhibit some level of loss at room temperature.
The one-port SRM standards were implemented as non-ideal match, short, and open standards with a $200\,\mu\mathrm{m}$ offset, as shown in Fig. 4. For the network-load standards, a $4\,\mathrm{mm}$ line is used as the reciprocal standard, combined with the non-ideal match, short, and open standards. Note that the CPW has a complex-valued, frequency-dependent characteristic impedance that deviates slightly from $50\,\Omega$. Thus, the CPW line segments introduce both an offset and impedance mismatch, as $Z_{\mathrm{ref}}=50\,\Omega$ is used as the reference impedance for the lumped elements. The device under test (DUT) used for verification is a stepped-impedance line using the same CPW structure but with a $15\,\mu\mathrm{m}$ signal width.
Figure 4: Models used to simulate non-ideal load standards (a) $50\,\Omega$ match standard with $L_{0}=25\,\mathrm{pH},C_{0}=1\,\mathrm{fF}$, (b) short standard with $L_{0}=30\,\mathrm{pH},L_{1}=10\times 10^{-12}\,\mathrm{pH/Hz},C_{0}=0.5\,\mathrm{fF}$, and (c) open standard with $C_{0}=15\,\mathrm{fF},C_{1}=5\times 10^{-15}\,\mathrm{fF/Hz},L_{0}=1\,\mathrm{pH}$. All standards include a $200\,\mu\mathrm{m}$ CPW line offset. These models are used for automatic parasitic extraction.
Only the match and short standards have been used for automatic model parameter estimation. All lumped elements and material properties ($\epsilon_{r}^{\prime}$, $\tan\delta$, and $\sigma_{r}$) were assumed unknown. While including the open standard would have been possible, it would have increased the parameter space from 11 parameters (five from the match standard and six from the short standard) to 17 parameters, significantly increasing optimization time. Symmetric models at both ports were assumed to improve optimization efficiency, as asymmetric models would double the parameter count. This consideration is particularly important as the CPW model, though analytical, relies on computationally intensive Bessel functions. In the measurement example in Section IV-B, devices have been modeled using only lumped elements without offset, allowing for more parameters and independent port solutions due to faster evaluation.
@@
To validate the accuracy of estimated parameters from the optimizer, a relative error metric has been defined:
@@
where $x_{\mathrm{est}}$ represents the estimated quantity and $x_{\mathrm{true}}$ is the true value. The estimated quantity corresponds to the parameter value retrieved from the optimization, while the true value is the actual value used to generate the simulation data. The relative error provides a normalized error metric that accommodates the different units of the different parameters.
Data processing was implemented in Python using the scikit-rf [31] and scipy [32] packages. The optimization was performed using the DE algorithm [27, 32], chosen for its robustness in handling high-dimensional parameter spaces without requiring gradient information or precise initial guesses. Furthermore, the SciPy implementation of DE supports parallelization, significantly accelerating convergence on multi-core processors compared to other global optimization routines available in the library.
Fig. 5 demonstrates the relative error of estimated parameters of the models used in SRM calibration, converging to floating-point precision. It should be noted that due to the computationally expensive CPW model, the 1000 iterations shown in Fig. 5 took approximately 25 minutes to complete (using 150 frequency points).
@@
To compare S-parameters, absolute error is used to avoid division by zero, particularly for reflection parameters. The error is defined in dB as follows:
$\text{Error}_{ij}(\mathrm{dB})=20\,\mathrm{log}_{10}\left|{S_{ij}^{\mathrm{SRM}}-S_{ij}^{\mathrm{Ref}}}\right|$ (14)
where $S_{ij}^{\mathrm{SRM}}$ denotes the S-parameter calibrated using the SRM method, and $S_{ij}^{\mathrm{Ref}}$ is the reference S-parameter. In this example, the reference corresponds to the actual simulated DUT before the error boxes are embedded. Fig. 6 shows the calibrated DUT results comparing two cases: using an ideal match standard versus using nonlinear optimization to fit unknown parameters. This ideal match models the match response by a frequency-independent $50\,\Omega$ resistor, even if the match for the simulation is calculated based on the model introduced in Fig. 4a. The assumption of frequency independence of the match is well applicable for frequencies below $30\,\mathrm{GHz}$. Above this frequency, the ideal match standard definition produces significant errors in both $S_{11}$ and $S_{21}$. The optimization approach achieves much better accuracy, limited only by numerical precision.
Figure 6: Comparison of the true S-parameters for the stepped impedance line (the actual simulated DUT prior to embedding with error boxes) against SRM calibration results obtained using the ideal match definition versus the model fitting procedure.
@@ IV-B PCB Measurement
For the multiline TRL standards, line lengths of $\{0,0.5,4,5.5,6.5,8.5\}\,\mathrm{mm}$ were employed, with the shortest line serving as the thru standard. The reflect standard was implemented as an open circuit, achieved by separating the probes in air. The multiline TRL algorithm used to process the measurements is based on the method described in [6, 7], which is als.
For SRM standards, symmetric open, short, and match standards, and network-load with an offset line segment were implemented. The match standard was implemented using a $50\,\Omega$ flip-chip Vishay CH0402-50RGFPA [35], specified for operation up to $50\,\mathrm{GHz}$ with a $\pm 2\%$ DC resistance tolerance. The DC resistance of the soldered flip-chip resistor was measured, indicating a value of $49\,\Omega$. This measurement was conducted by first zeroing the ohmmeter, which was connected through the bias-tee behind the probe at Port-A, and measuring a short standard implemented as a microvia. Subsequently, the soldered flip-chip resistor at Port-A was measured. The same measured DC value was used for port-B. The reference impedance of the SRM calibration was set to a constant reference impedance of $50\,\Omega$. It should be noted that due to the small size of the resistors (0402 footprint), manual soldering was challenging, potentially introducing asymmetry in frequency response due to the solder bumps. Fig. 7 illustrates the measurement system and the soldered resistors.
@@
For comparison between SRM and multiline TRL, a stepped impedance microstrip line was used as DUT with $100\,\mu\mathrm{m}$ strip width. For proper comparison, the multiline TRL calibration was renormalized to a constant reference impedance of $50\,\Omega$. Fig. 8 shows the fitting models used for short, open, and match standards in SRM calibration. Both open and short standards were included in the optimization since they only comprise lumped elements and their evaluation is computationally efficient. The fitting was performed independently at each port. The total optimization time for both ports was approximately four minutes, fitting 22 parameters (using 197 frequency points). Only the parameters associated with the match results are used in the final SRM calibration. The resistor model follows the datasheet specifications [35], with the microvia termination modeled as an inductor.
@@
Fig. 9 compares the S-parameters of the calibrated DUT between multiline TRL and SRM. The DUT response from multiline TRL calibration was used as the reference for calculating the absolute error based on (14). For SRM calibration, several scenarios were examined for defining the match standard: First, assuming ideal zero reflection (typical choice when frequency-dependent response is unknown); second, using our optimization procedure from Section III to fit the models in Fig. 8; and finally, using the measurement of the match standard based on the multiline TRL calibration.
Results in Fig. 9 demonstrate that assuming an ideal match response leads to substantial errors, particularly above $10\,\mathrm{GHz}$ in both $S_{11}$ and $S_{21}$. In contrast, the optimization approach yields errors comparable to defining the match standard based on multiline TRL. Some discrepancies exist between the DUT response based on SRM and multiline TRL calibrations, especially near $50\,\mathrm{GHz}$. This deviation may arise because the match standard exhibits behavior similar to a short standard near $50\,\mathrm{GHz}$, as evident in Fig. 10, thereby increasing the sensitivity of SRM calibration due to insufficient unique standards and causing the calibration problem to be ill-conditioned. Additionally, the discrepancy could stem from inconsistent resistor soldering, as asymmetry in the standards affects the calibration as investigated in [20]—an error that becomes amplified when the match standard behaves like a short standard near $50\,\mathrm{GHz}$.
Nevertheless, achieving results comparable to multiline TRL, with an absolute error below $22\,\mathrm{dB}$ for frequencies under $20\,\mathrm{GHz}$, is notable given the use of manually soldered flip-chip resistors on a PCB with significant parasitic responses. This is particularly significant considering only the DC resistance was specified during the SRM calibration. In contrast, assuming an ideal match yields poor results, exceeding $10\,\mathrm{dB}$ at $20\,\mathrm{GHz}$ with significant degradation at higher frequencies. Furthermore, even for multiline TRL, the expected impedance variation of the traces can range from $\pm 4\,\Omega$ [21]; hence, deviations between SRM and multiline TRL at lower frequencies are expected due to inherent impedance variations in PCB traces.
It is also worth noting that the SRM method with an optimized model yields a response similar to that obtained with SRM calibration and a defined match from multiline TRL measurements. This indicates that the deviation lies not in the definition of the match standard for the SRM method, but rather in the discrepancy between the standards. At higher frequencies, specifically above $20\,\mathrm{GHz}$, the match standard becomes increasingly reflective, as depicted in Fig. 10. Better agreement with multiline TRL at higher frequencies could be achieved by reducing resistor parasitics—for example, by using machine-placed resistors or higher-quality resistors where pad parasitics are less influential—or alternatively, by using multiple offset-short standards with different lengths to maintain a stable calibration problem at higher frequencies, as implemented in some coaxial calibration kits [36].